# NICER Magnetar Burst Catalog

### v1

@@ Abstract
In this paper, we present a comprehensive catalog of short bursts from magnetars based on eight years of NICER observations. A total of 1130 bursts were identified, making this the largest magnetar burst catalog to date. The sample is dominated by SGR 1935+2154, which contributes 76% of all detected bursts. We analyzed burst durations, spectral properties, and their correlations across multiple sources. Bursts from SGR 1935+2154 exhibit significantly longer durations, with a mean of 317 ms, compared to a mean of 23 ms for bursts from other magnetars. Two s-scale bursts were detected for the first time, originating from 1E 1048.15937 and CXOU J010043.1721134. Spectral analysis in the 0.5–8 keV range using both blackbody and power-law models shows that bursts with higher fluences have harder spectra. In contrast, correlations between burst duration and spectral parameters are weak or absent. This catalog provides a valuable dataset for studying magnetar short bursts, enabling future modeling efforts and improving our understanding of the diversity and physical mechanisms of magnetar bursts.
1 Introduction
Magnetars are a class of neutron stars characterized by their extremely strong magnetic fields, often exceeding – G (1992ApJ...392L...9D). In the X-ray band, unlike rotation-powered pulsars, which typically show steady emission, magnetars are known for their highly variable and transient behavior (e.g., 2018MNRAS.474..961C). The first discovered magnetar burst was the giant flare from SGR 052666 in 1979, followed by a short burst (1979Natur.282..587M). Shortly afterwards, a series of short bursts was detected from SGR 1900+14 (1979SvAL....5..343M). Because these repeating events are dominated by emission in the soft -ray band, the sources were named soft gamma repeaters (SGRs). During a magnetar outburst, which can last from several months to several years, short bursts often appear in clusters known as burst forests or burst storms, with rates that can reach hundreds of bursts per hour (e.g., 2020ApJ...904L..21Y; 2025ApJ...989...63H). In contrast, during the quiescent state, bursts may appear individually (e.g., 2002Natur.419..142G). According to the magnetar model, these bursts are generally thought to be triggered by cracks in the neutron star crust, induced by extreme magnetic stress (1992ApJ...392L...9D; 1995MNRAS.275..255T).
@@ 1 Introduction
A more comprehensive burst catalog was later compiled using data from Fermi/Gamma-ray Burst Monitor (GBM), which included bursts from more than eight magnetar sources, though the majority originated from 1E 1547.05408 (2015ApJS..218...11C). The timing and spectral properties reported in this catalog are consistent with those obtained from earlier observations. In the 2020s, several magnetars underwent new outbursts that produced numerous short bursts. Detailed analyses have been conducted for bursts from Swift J1555.25402, SGR 18300645, and SGR 1935+2154 (2021ApJ...920L...4E; 2022ApJ...924..136Y; 2025ApJ...989...63H). While their general properties are similar to those of other magnetar bursts, the events from SGR 1935+2154 stand out with substantially longer durations, peaking around 1 s, which is significantly longer than other magnetars (2020ApJ...904L..21Y).
@@
2 Data and Analysis
2.1 Data Selection and Reduction
NICER is a soft X-ray (0.2–12 keV) telescope mounted on the International Space Station (2016SPIE.9905E..1HG). Its X-ray Timing Instrument (XTI) consists of 56 Focal Plane Modules (FPMs), though only 52 modules were in operation. While it does not offer imaging capability, NICER provides exceptionally high time resolution ( 100 ns) and a large effective area, making it ideal for studying transient events. For this work, we selected all NICER observations conducted from its launch, 2017 June, through 2025 June 30 that included a magnetar or a magnetar-like pulsar within the field of view. Among the 26 known magnetars (2014ApJS..212....6O)111McGill Online Magnetar Catalog http://www.physics.mcgill.ca/~pulsar/magnetar/main.html, NICER has observed 21 of them, along with two additional magnetar-like pulsars. We collected 6016.9 ks of exposure time from 2765 NICER observations across 23 sources. The total exposure time and number of observations used for each source are summarized in Table 1.
@@ 2.1 Data Selection and Reduction
2.2 Burst Searching
We applied the Bayesian block algorithm to both the standard and high-energy event lists to identify significant flux variations (1998ApJ...504..405S). Blocks occurring during high-particle-background intervals, as defined by the high-energy band, were excluded. In the standard band, blocks with durations shorter than 1 s were selected as burst candidates. We further examined the FPM distribution of each candidate. Candidates with events concentrated in only one or two FPMs were identified as particle-induced events and excluded from the burst sample.
@@ 2.2 Burst Searching
NICER has been monitoring seven X-ray blank regions defined initially by the Rossi X-ray Timing Explorer (RXTE; 2006ApJS..163..401J). We selected one observation from each region and excluded intervals with high particle background. The details of the blank sky observations used are summarized in Table 2. We binned the cleaned event lists into 0.2 s and 0.02 s bins, matching the burst durations reported in previous work (2015ApJS..218...11C) as well as those found in this study. We then recorded the total counts in each bin to characterize the background X-ray distribution. The cumulative distribution function reaches 100% at 4 counts per 0.2 s bin and 3 counts per 0.02 s bin (Figure 1). To ensure robust separation between genuine bursts and background fluctuations, we required that a burst candidate contain at least 5 counts within a Bayesian block. Based on the measured background rate and the NICER effective area, the expected detection limits for a transient event are approximately 2 and 2.5 erg s-1 cm-2 for transient duration 0.02 s and 0.2 s respectively.
@@ 2.3 Analysis
where soft denotes photons in the 0.5–4 keV range and hard denotes photons in the 4–8 keV range. Some bursts show no photons below 2 keV because of strong absorption along the line of sight, so we adopted 4 keV as the boundary between the soft and hard photons, rather than 2 keV.
@@
3 Results and Discussion
3.1 Overall samples
@@
Several magentar outbursts have also been captured by NICER since its launch in 2017 June, including the 2020 outburst of Swift J1818.01607 (2020ApJ...902....1H), the 2020 outburst of SGR 18300645 (2022ApJ...924..136Y), the 2021 outburst of Swift J1555.25402 (2021ApJ...920L...4E), and the 2024 outburst of 1E 1841045 (2025ApJ...989...89Y). These sources accounted for 21.6% of the bursts in our sample. NICER also observed additional outbursts from other magnetars, such as XTE J1810197 (2019ApJ...877L..30G), and PSR J18460258 (2023ApJ...952..120H). However, in these cases, the outburst activity was relatively weak or the observations did not coincide with the outburst onset, resulting in fewer detected bursts.
@@
3.2 duration
The distribution of durations for all bursts is shown in Figure 2. The overall sample has a mean of 350 ms and a standard deviation of 470 ms. Two s-scale bursts were identified in our sample: one from 1E 1048.15937 with a duration of 0.6 s, and another from CXOU J010043.1721134 with a duration of 5 s. These represent the first detections of s-scale bursts in X-rays from magnetars. Their light curves are shown in Figure 3. Since no sub-ms bursts were found, it remains unclear whether these events are simply the extreme tail of the main burst distribution or instead represent a distinct class of magnetar short bursts. Future X-ray missions with larger effective area will be crucial to resolve this question.
Bursts from SGR 1935+2154 exhibit values that are significantly longer than those of other magnetars. We fitted log-normal distributions to both groups and show the normalized distributions in Figure 2. For SGR 1935+2154, the distribution peaks at 316 ms, while it peaks at 23 ms for the remaining magnetars. Thus, bursts from SGR 1935+2154 are on average about an order of magnitude longer. A Kolmogorov–Smirnov (KS) test applied to the two distributions yields a probability of , confirming that the two samples are distinct.
@@
2020ApJ...904L..21Y studied bursts from SGR 1935+2154 during its 2020 outburst using NICER data. These events are also included in our sample, accounting for roughly 40% of the SGR 1935+2154 bursts analyzed here. However, the durations we obtain are systematically shorter than those reported in 2020ApJ...904L..21Y. A key reason is the different treatment of multi-peak bursts: we apply a 0.2 s criterion to decide whether peaks belong to separate bursts, whereas their study used 0.5 s. Consequently, they classified more events as multi-peak bursts, yielding shorter durations if treated as individual bursts. A second factor is the difference in burst populations across epochs. NICER captured a burst storm from SGR 1935+2154 in the 2020 outburst, during which the persistent emission was higher than usual. Dim or very short bursts may have been buried in this elevated background and thus missed. In contrast, the bursts detected in the 2021 and 2022 outbursts occurred during lower persistent-emission states, making weaker and shorter events easier to identify. As a result, the 2021 and 2022 bursts exhibit shorter duration distributions overall, pulling the combined distribution toward smaller values.
@@
In addition to bursts from other magnetars, 2013ApJ...778..105L reported 275 bursts from 1E 1547.05048 detected with Swift/X-ray Telescope (XRT), with a distribution peaking at 207 ms and 262 bursts from SGR 0501+4516 detected with XMM-Newton, whose durations peak at about 100 ms. 2015ApJS..218...11C presented a catalog of 440 bursts from at least eight magnetars observed by Fermi/GBM. Among these, 386 bursts originate from 1E 1547.05048 and have a peak at 156 ms, while the remaining bursts peak around 80 ms. Overall, these duration distributions are longer than those of the regular magnetar bursts in our NICER sample. Because NICER offers superior time resolution and larger effective area than those aforementioned missions, it is more sensitive to faint or shorter duration bursts. As a result, our catalog contains many more short events, leading to systematically shorter distributions compared to previous studies. A summary of the log-normal fitting results of the distributions, including the peak values and 16–84% quantile ranges for different groups and instruments, is presented in Table 4.
3.3 Hardness ratio
@@ 3.4 Spectral model
The results of the blackbody fits are shown in Figure 5, which presents the distributions of blackbody temperature () and flux (). Fitting the distribution of SGR 1935+2154 group with a log-normal function yields a peak at 1.56 keV and a 16–84% quantile range of 1.17–2.08 keV. This characteristic temperature is slightly lower than that obtained from Swift/XRT bursts of 1E 1547.05408 (2.5 keV; 2013ApJ...778..105L) and considerably lower than that of Fermi/GBM bursts (4.5 keV; 2015ApJS..218...11C). The difference likely reflects NICER’s higher sensitivity to fainter bursts, which tend to exhibit softer spectra (2001ApJ...558..228G; 2013ApJ...778..105L), corresponding to lower blackbody temperatures (see also Section 3.5). As a result, the inclusion of additional dim bursts shifts the overall distribution toward lower values. The higher temperatures inferred from Fermi/GBM may also arise because those fits correspond to the lower-temperature component of a two-blackbody model in the 8–200 keV range, where the cooler component is poorly constrained and thus biased upward. Alternatively, the broadband spectra may be better represented by three blackbody components rather than two. In any case, the burst temperatures are significantly higher than the typical magnetar temperature in quiescence of 0.45 keV (2023MNRAS.526.1287C), and even during the onset of outburst of 0.8 keV (e.g. 2019ApJ...874L..25G). The fluxes are generally distributed between and erg s-1 cm-2, with the two s-scale bursts showing much higher fluxes of erg s-1 cm-2.
@@
From the turnover in these distributions, we estimate that NICER achieves nearly 100% detection efficiency above erg cm-2, with a detection limit of approximately erg cm-2. For the duration distribution of our burst sample, this detection limit is slightly higher than the expected value of erg cm-2 estimated in Section 2.2. The reason is that most of the bursts identified in this work occurred during a magnetar X-ray outburst, where the persistent emission is higher. As a result, the background level is higher, and some low-fluence bursts are buried in the background.
3.5 Correlation
The left panels of Figure 8 show the distributions of and versus their corresponding fluxes. For low-count bursts, where or were fixed during fitting, these points are not included in the figure. Overall, bursts with higher fluxes tend to exhibit harder spectra, characterized by higher temperatures or smaller photon indices, consistent with previous studies (see, e.g., 2001ApJ...558..228G; 2013ApJ...778..105L). To quantify this relation, we calculated Pearson correlation coefficients (Pearson r) for bursts from SGR 1935+2154 and from other magnetars separately. The clear outlier, the 5 s burst from CXOU J010043.1721134, was excluded from this analysis.

### v2

@@ Abstract
In this paper, we present a comprehensive catalog of short bursts from magnetars based on eight years of NICER observations. A total of 1130 bursts were identified and the sample is dominated by SGR 1935+2154, which contributes 76% of all detected bursts. We analyzed burst durations, spectral properties, and their correlations across multiple sources. Bursts from SGR 1935+2154 exhibit significantly longer durations, with a mean of 317 ms, compared to a mean of 23 ms for bursts from other magnetars. Two s-scale bursts were detected for the first time, originating from 1E 1048.15937 and CXOU J010043.1721134. Spectral analysis in the 0.5–8 keV range using both blackbody and power-law models shows that bursts with higher fluences have harder spectra. In contrast, correlations between burst duration and spectral parameters are weak or absent. This catalog provides a valuable dataset for studying magnetar short bursts, enabling future modeling efforts and improving our understanding of the diversity and physical mechanisms of magnetar bursts.
I Introduction
Magnetars are a class of neutron stars characterized by their extremely strong magnetic fields, often on the order of – G (1992herm.book.....M; 1992ApJ...392L...9D; 1992AcA....42..145P). In the X-ray band, unlike rotation-powered pulsars, which typically show steady emission, magnetars are known for their highly variable and transient behavior (e.g., 2018MNRAS.474..961C). The first discovered magnetar burst was the giant flare from SGR 052666 in 1979, with a duration of about 10 minutes (1979Natur.282..587M). Shortly afterwards, a series of short bursts, with durations ranging from milliseconds to seconds, was detected from SGR 1900+14 (1979SvAL....5..343M). Because these repeating events are dominated by emission in the soft -ray band, the sources were named soft gamma repeaters (SGRs). During a magnetar outburst, which can last from several months to several years, short bursts often appear in clusters known as burst forests or burst storms, with rates that can reach hundreds of bursts per hour (e.g., 2020ApJ...904L..21Y; 2025ApJ...989...63H). In contrast, during the quiescent state, bursts may appear individually (e.g., 2002Natur.419..142G). According to the magnetar model, these bursts are generally thought to be triggered by activity in the neutron star crust, induced by extreme magnetic stresses (1984Natur.310..121L; 1992ApJ...392L...9D; 1995MNRAS.275..255T).
@@ I Introduction
A more comprehensive burst catalog was later compiled using data from Fermi/Gamma-ray Burst Monitor (GBM), which included bursts from more than eight magnetar sources, with the majority originated from 1E 1547.05408 (2015ApJS..218...11C). More recently, the INTErnational Gamma-Ray Astrophysics Laboratory (INTEGRAL)/Imager on Board the INTEGRAL Satellite (IBIS) combined over 20 years of observations to construct a catalog of 1349 magnetar bursts (2025arXiv251214356P). The timing and spectral properties reported in this catalog are consistent with those obtained from earlier observations. In the 2020s, several magnetars underwent new outbursts that produced numerous short bursts. Detailed analyses have been conducted for bursts from Swift J1555.25402, SGR 18300645, and SGR 1935+2154 (2021ApJ...920L...4E; 2022ApJ...924..136Y; 2025ApJ...989...63H). While their general properties are similar to those of other magnetar bursts, the events from SGR 1935+2154 stand out with substantially longer durations, peaking around 1 s, which is significantly longer than other magnetars (2020ApJ...904L..21Y).
@@
II Data and Analysis
II.1 Data Selection and Reduction
NICER is a soft X-ray (0.2–12 keV) telescope mounted on the International Space Station (2016SPIE.9905E..1HG). Its X-ray Timing Instrument (XTI) consists of 56 Focal Plane Modules (FPMs), though only 52 modules were in operation. While it does not offer imaging capability, NICER provides exceptionally high time resolution ( 300 ns) and a large effective area, making it ideal for studying transient events. For this work, we selected all NICER observations conducted from its launch, 2017 June, through 2025 June 30 that included a magnetar or a magnetar-like pulsar within the field of view. Among the 26 known magnetars (2014ApJS..212....6O)111McGill Online Magnetar Catalog http://www.physics.mcgill.ca/~pulsar/magnetar/main.html, NICER has observed 21 of them, along with two additional magnetar-like pulsars. We collected 6016.9 ks of exposure time from 2765 NICER observations across 23 sources. The total exposure time and number of observations used for each source are summarized in Table 1.
@@ II.1 Data Selection and Reduction
II.2 Burst Searching
We applied the Bayesian block algorithm implemented in the astropy package, using the fitness=events option, to both the standard and high-energy event lists to identify significant flux variations (1998ApJ...504..405S; 2013ApJ...764..167S). Blocks occurring during high-particle-background intervals, as defined by the high-energy band, were excluded. In the standard band, blocks with durations shorter than 1 s were selected as burst candidates. We further examined the FPM distribution of each candidate. Candidates with events concentrated in only one or two FPMs were identified as particle-induced events and excluded from the burst sample.
@@ II.2 Burst Searching
NICER has been monitoring seven X-ray blank regions defined initially by the Rossi X-ray Timing Explorer (RXTE; 2006ApJS..163..401J). We selected one observation from each region and excluded intervals with high particle background. The details of the blank sky observations used are summarized in Table 2. The RXTE7 region is excluded because it was later found to contain a faint X-ray source (2022AJ....163..130R). In addition, some of the remaining regions contain fainter sources (up to 3 erg s-1 cm-2) according to the extended ROentgen Survey with an Imaging Telescope Array (eROSITA) source catalog (2024A&A...682A..34M). We binned the cleaned event lists into 0.2 s and 0.02 s bins, matching the burst durations reported in previous work (2015ApJS..218...11C) as well as those found in this study. We then recorded the total counts in each bin to characterize the background X-ray distribution. The cumulative distribution function reaches 100% at 4 counts per 0.2 s bin and 3 counts per 0.02 s bin (Figure 1). To ensure robust separation between genuine bursts and background fluctuations, we required that a burst candidate contain at least 5 counts within a Bayesian block. Based on the measured background rate and the NICER effective area, the expected detection limits for a transient event are approximately 2 and 2.5 erg cm-2 for transient duration 0.02 s and 0.2 s respectively.
@@ II.3 Analysis
where soft denotes photons in the 0.5–4 keV range and hard denotes photons in the 4–8 keV range. Some bursts show no photons below 2 keV because of strong absorption along the line of sight, so we adopted 4 keV as the boundary between the soft and hard photons, rather than the more typical 2 keV value.
@@
III Results and Discussion
III.1 Overall samples
@@
Several magnetar outbursts have also been captured by NICER since its launch in 2017 June, including the 2020 outburst of Swift J1818.01607 (2020ApJ...902....1H), the 2020 outburst of SGR 18300645 (2022ApJ...924..136Y), the 2021 outburst of Swift J1555.25402 (2021ApJ...920L...4E), and the 2024 outburst of 1E 1841045 (2025ApJ...989...89Y). These sources accounted for 21.6% of the bursts in our sample. NICER also observed additional outbursts from other magnetars, such as XTE J1810197 (2019ApJ...877L..30G), and PSR J18460258 (2023ApJ...952..120H). However, in these cases, the bursting activity was relatively weak and/or the observations did not coincide with the outburst onset, when the strongest bursting activity is typically observed.
@@
III.2 duration
The distribution of durations for all bursts is shown in Figure 2. The overall sample has a mean of 350 ms and a standard deviation of 470 ms. Two s-scale bursts were identified in our sample: one from 1E 1048.15937 with a duration of 0.6 s, and another from CXOU J010043.1721134 with a duration of 5 s. We verified that these s-scale bursts are not associated with known instrumental effects. The absolute timing precision of NICER ( 300 ns) is well below the measured burst durations. In each burst, the detected photons are distributed across multiple FPMs, with no single FPM registering more than one hit, which makes an instrumental or electronic origin unlikely. These detections represent the first observations of s-scale X-ray bursts from magnetars. Their light curves are shown in Figure 3. Since no bursts were detected with durations between 10 s and 1 ms, it remains unclear whether these events are simply the extreme tail of the main burst distribution or instead represent a distinct class of magnetar short bursts. Future X-ray missions with larger effective area will be crucial to resolve this question.
Bursts from SGR 1935+2154 exhibit values that are significantly longer than those of other magnetars. We fitted log-normal distributions to both groups and show the normalized distributions in the right panel of Figure 2. For SGR 1935+2154, the distribution peaks at 316 ms, while it peaks at 23 ms for the remaining magnetars. Thus, bursts from SGR 1935+2154 are on average about an order of magnitude longer. A Kolmogorov–Smirnov (KS) test applied to the two distributions yields a probability of , confirming that the two samples are distinct. Nevertheless, a small subset of bursts from other magnetars extends to longer durations comparable to those of SGR 1935+2154, suggesting that long-duration bursts are not unique to this source and may be associated with source-dependent activity states.
@@
2020ApJ...904L..21Y studied bursts from SGR 1935+2154 during its 2020 outburst using NICER data. These events are also included in our sample, accounting for roughly 40% of the SGR 1935+2154 bursts analyzed here. However, the durations we obtain are systematically shorter than those reported in 2020ApJ...904L..21Y. A key reason is the different treatment of multi-peak bursts: we apply a 0.2 s criterion to decide whether peaks belong to separate bursts, whereas their study used 0.5 s. Consequently, they grouped more peaks into single multi-peak bursts, resulting in longer measured durations. A second factor is the difference in burst populations across epochs. NICER captured a burst storm from SGR 1935+2154 in the 2020 outburst, during which the persistent emission was higher than usual. Dim or very short bursts may have been buried in this elevated background and thus missed. In contrast, the bursts detected in the 2021 and 2022 outbursts occurred during lower persistent-emission states, making weaker and shorter events easier to identify. As a result, the 2021 and 2022 bursts exhibit shorter duration distributions overall, pulling the combined distribution toward smaller values.
@@
In addition to bursts from other magnetars, 2013ApJ...778..105L reported 275 bursts from 1E 1547.05048 detected with Swift/X-ray Telescope (XRT), with a distribution peaking at 207 ms and 262 bursts from SGR 0501+4516 detected with XMM-Newton, whose durations peak at about 100 ms. 2015ApJS..218...11C presented a catalog of 440 bursts from at least eight magnetars observed by Fermi/GBM. Among these, 386 bursts originate from 1E 1547.05048 and have a peak at 156 ms, while the remaining bursts peak around 80 ms. Overall, these duration distributions are longer than those of the regular magnetar bursts in our NICER sample. One likely explanation is that NICER offers superior time resolution and larger effective area than those aforementioned missions, it is more sensitive to faint or shorter duration bursts. Consequently, our catalog includes a substantial number of short bursts, leading to systematically shorter distributions compared to previous studies.
Differences in energy coverage and background levels also play an important role. The Fermi/GBM catalog is based on observations in the 8–200 keV range, which differs significantly from the 0.5–8 keV band used in this work. In addition, Fermi/GBM has a higher background rate, which can cause the rising and decaying tails of bursts to be buried in the background. As a result, burst durations measured with GBM are often shorter than those measured with NICER, even for the same burst, as shown by 2021NatAs...5..408Y. This effect is also evident for 1E 1547.05048, where the peak is longer in Swift/XRT data than in Fermi/GBM data, despite probing the same outburst (2013ApJ...778..105L; 2015ApJS..218...11C). In contrast, most bursts from magnetars other than SGR 1935+2154 in our NICER catalog (e.g., Swift J1555.25402, Swift J1818.01607, and SGR 18300645) exhibit durations significantly shorter than those of 1E 1547.05048 detected by Fermi/GBM. This difference suggests that the bursts captured by GBM may be primarily sourced from more energetic bursting episodes, similar to the highly active 2020 outburst of SGR 1935+2154. Interestingly, the 2009 outburst of 1E 1547.05048, from which the GBM bursts originate, was also accompanied by the detection of radio bursts (2021ApJ...907....7I). Although these radio bursts were less energetic than typical extragalactic FRBs and the FRBs associated with SGR 1935+2154, their occurrence may point to the existence of different magnetar outburst modes, some of which are characterized by more energetic X-ray bursting activity and an enhanced likelihood of radio burst emission.
A summary of the log-normal fitting results of the distributions, including the peak values and 16–84% quantile ranges for different groups and instruments, is presented in Table 4.
III.3 Hardness ratio
@@ III.4 Spectral model
The results of the blackbody fits are shown in Figure 5, which presents the distributions of blackbody temperature () and flux (). Fitting the distribution of SGR 1935+2154 group with a log-normal function yields a peak at 1.56 keV and a 16–84% quantile range of 1.17–2.08 keV. This characteristic temperature is slightly lower than that obtained from Swift/XRT bursts of 1E 1547.05408 (2.5 keV; 2013ApJ...778..105L) and considerably lower than that of Fermi/GBM bursts (4.5 keV; 2015ApJS..218...11C). The difference likely reflects NICER’s higher sensitivity to fainter bursts, which tend to exhibit softer spectra (2001ApJ...558..228G; 2013ApJ...778..105L), corresponding to lower blackbody temperatures (see also Section III.5). As a result, the inclusion of additional dim bursts shifts the overall distribution toward lower values. The higher temperatures inferred from Fermi/GBM may also arise because those fits correspond to the lower-temperature component of a two-blackbody model in the 8–200 keV range, where the cooler component is poorly constrained and thus biased upward. Alternatively, the broadband spectra may be better represented by three blackbody components rather than two. In any case, the burst temperatures are significantly higher than the typical magnetar temperature in quiescence of 0.45 keV (2023MNRAS.526.1287C), and even during the onset of outburst of 1.2 keV (e.g. 2020ApJ...902....1H; 2021ApJ...920L...4E). The fluxes are generally distributed between and erg s-1 cm-2, with the two s-scale bursts showing much higher fluxes of erg s-1 cm-2.
@@
From the turnover in these distributions, we estimate that NICER achieves nearly 100% detection efficiency above erg cm-2, with a detection limit of approximately erg cm-2. For the duration distribution of our burst sample, this detection limit is slightly higher than the expected value of erg cm-2 estimated in Section II.2. The reason is that most of the bursts identified in this work occurred during a magnetar X-ray outburst, where the persistent emission is higher. As a result, the background level is higher, and some low-fluence bursts are buried in the background.
III.5 Correlation
The left panels of Figure 8 show the distributions of and versus their corresponding fluxes. For low-count bursts, where or were fixed during fitting, these events are excluded from the figure and from the correlation analysis. Overall, bursts with higher fluxes tend to exhibit harder spectra, characterized by higher temperatures or smaller photon indices, consistent with previous studies (see, e.g., 2001ApJ...558..228G; 2013ApJ...778..105L). To quantify this relation, we calculated Pearson correlation coefficients (Pearson r) for bursts from SGR 1935+2154 and from other magnetars separately. The clear outlier, the 5 s burst from CXOU J010043.1721134, was excluded from this analysis.